OpenStudy (anonymous):

7. A ball is dropped vertically from a height of 50 meters. After each bounce, the ball reaches a maximum height equal to 80 percent of the maximum height of the previous bounce. a.) Write the first three terms of the sequence. Explain how you know what they are. b.) Find the height to the nearest tenth of a meter of the ball after the tenth bounce. Show and explain your work.

OpenStudy (anonymous):

can someone help me?(:

OpenStudy (anonymous):

lol... really...

OpenStudy (anonymous):

this isnt multiple choice. just a 2 part physics problem

OpenStudy (whpalmer4):

After the first bounce, the ball reaches a height of 50*0.8 = 40 meters After the second bounce, the ball reaches a height of 40*0.8 = 32 meters. Can you work out the third bounce height?

OpenStudy (anonymous):

im confused on how to solve them, but i think a would be .8(50) = 40m .64 (50) = 32m .512 (50) = 25.6m am i right?

OpenStudy (whpalmer4):

After \(n\) bounces, the height is \(0.8^{n}*50\)

OpenStudy (whpalmer4):

Yes, those are correct

OpenStudy (anonymous):

0.8^5(50) = 16.38?

OpenStudy (whpalmer4):

Aren't you supposed to do the 19th bounce? And round to the nearest 10th of a meter?

OpenStudy (whpalmer4):

10th bounce, missed the key, it appears!

OpenStudy (anonymous):

so .8 ^10 (50) = 5.4 ?